OpenStudy (anonymous):

I thought I was doing it right until I went to mark my answer and the answer I got was not there, so please help me question 9squ7-squ28 I got for my answer 5qu7??

OpenStudy (anonymous):

@phi

OpenStudy (phi):

\[ 9 \sqrt{7} - \sqrt{28} \] factor 28 into its prime factors, and pull out pairs 28= 2*14= 2*2*7 so \[ \sqrt{28}= \sqrt{2\cdot 2 \cdot 7} \] there is one pair of 2's we can take out (and put only 1 of them out front) we get \[ \sqrt{2\cdot 2 \cdot 7} = 2 \sqrt{7}\] now your problem is \[ 9 \sqrt{7} - 2\sqrt{7} \]

OpenStudy (phi):

if you have 9 sqr(7) and take away 2 of them, how many sqr(7)'s do you have ?

OpenStudy (anonymous):

it would be 7squ7 thank you I don't know what I wronge I have another that im stuck on too

OpenStudy (phi):

please post it.